OpenStudy (anonymous):

f(w) = w^4+3w^3-22w^2 + 2 is f'(w) = 4w^3 +9w^2 -44w right? I want to set this equal to 0, so can i do this? 4w^3 +9w^2 -44w = 0 w(4w^2 +9w -44) = 0 and then use a quadratic formula? if i do, what happen to the w at the left?

OpenStudy (ranga):

If A * B = 0 then either A = 0 or B = 0 or both are 0. A cubic equations will have 3 roots. w(4w^2 +9w -44) = 0 implies w = 0 or 4w^2 +9w -44 = 0 solving the quadratic will give you two roots. The third root is w = 0.

OpenStudy (anonymous):

\[-9 \pm \sqrt{9^2 - 4(4)(-44)} \] is equal to 1.9949 and - 4.2449 right?

OpenStudy (ranga):

Don't forget to divide by 2a or 8

OpenStudy (anonymous):

(−9±92−4(4)(−44))/2(4)

OpenStudy (ranga):

Above 9^2 looks like 92 and square root is missing. Use your reply before the last one but just divide by 8.

OpenStudy (ranga):

w = 1/8(-9 + sqrt(785)) or 1/8(-9 - sqrt(785)) and the third root is w = 0.

OpenStudy (anonymous):

Oh yes, when i copy and paste, the square has disappear